# Generalized symmetry in non-Hermitian systems

### v1

@@ I Introduction
Dynamics of non-Hermitian systems is a new frontier of exploration which harbors a lot of potential to discover new physical paradigms. A plethora of novel phenomena, such as dynamically induced topological phases, exponentially-enhanced quantum sensing [26] and the divergence of Lieb-Robinson bounds [8, 27] have been reported recently. Despite its increasing popularity, there is a lack of consensus in the mathematical formulation of non-Hermitian quantum mechanics in the community. As non-Hermitian time evolution is intrinsically non-unitary, different methods have been adopted to obtain a normalized physical observable in non-Hermitian systems. Among them is biorthogonal quantum mechanics (BQM) [28], which is a comprehensive formulation of non-Hermitian quantum mechanics consistent with its probabilistic interpretation. BQM is widely adopted within the condensed matter community [29, 30, 31] and mainly applied to study systems featuring a time-independent Hamiltonian. The so-called metric framework  [32, 33, 34] provides a generalization of BQM to the case with time-dependent Hamiltonians. In this framework, one considers a modified Hilbert space endowed by a time-dependent metric which appears as a weight factor in the inner product. The equation of motion of the metric incorporates the dynamics induced by both the Hamiltonian and its Hermitian conjugate. Both of these methods generate an exact time evolution of the non-Hermitian system, and are best suited when the non-Hermiticity of the system is realized in a fundamental manner instead of arising as an approximation. For example, this can be realized in NV centres via Naimark dilation [12].
On the other hand, a lot of the available literature on non-Hermitian dynamics [35, 36, 37, 38, 27, 39] skirts the important issue of non-conserved norms during time evolution and simply divides physical quantities by the time-dependent norm of the time-evolved state. This approach is ubiquitous in the open quantum systems community and is motivated by the quantum trajectory theory [40, 41, 42]. Quantum trajectories are used to find the time evolution of open quantum systems [43], for example when the direct solution of Lindblad master equation is computationally costly. In certain scenarios such as the detection of photons emitted from a cavity, the evolution of the system can be approximated by that with an effective non-Hermitian Hamiltonian interrupted by quantum jumps at random times [40]. By post-selecting quantum trajectories which did not undergo quantum jumps, a non-Hermitian evolution can be realized and the time-evolved state has to be explicitly normalized by hand at every time step [35].
@@ II.1 Biorthogonal quantum mechanics (BQM)
We start with a brief summary of BQM [28] for a general static non-Hermitian Hamiltonian $H$ with a non-degenerate spectrum. We assume a finite-dimensional Hilbert space throughout the paper. The inifinite-dimensional case has been discussed in Ref. [28]. The left and right eigenstates of $H$ are defined as
@@
as expectation values of observables in the biorthogonal framework are defined as $\langle\hat{F}\rangle=\tilde{\langle\psi\lvert}\hat{F}\lvert\psi\rangle$. In particular, setting $\hat{F}=\mathbbm{1}$, we see $\langle\tilde{\psi}\lvert\psi\rangle=\sum_{n}|c_{n}|^{2}$. We note that this does not necessarily equate to one, as $\sum_{n}\lvert n\rangle_{L~{}L}\langle n\lvert$ is not necessarily equal to the identity.
@@ II.2 Metric framework
Time evolution in the Hilbert spaces $\mathscr{H}_{\rho(t)}$   and $\mathscr{H}$  is generated by the respective Hamiltonians, $H(t)$ and $h(t)$, via the time-dependent Schrödinger equation (TDSE)
@@
where the states are related by $|\Psi(t)\rangle=\eta(t)|\psi(t)\rangle$. The construction of $\rho(t)$ guarantees that $h(t)$ is Hermitian, such that the unitarity of the time evolution is restored, giving $\langle\psi(t)|\rho(t)|\psi(t)\rangle=\langle\Psi(t)|\Psi(t)\rangle=1$ at all times $t$.
@@ II.3 Open quantum systems
and neglecting the term $\Gamma\rho_{\mathrm{me}}\Gamma^{\dagger}$, which is known as the recycling or quantum jump term. Curiously, the approximate master equation in the no-jump limit, $\dot{\rho}_{\mathrm{nj}}=-i(H_{\mathrm{eff}}\rho_{\mathrm{nj}}-\rho_{\mathrm{nj}}H_{\mathrm{eff}}^{\dagger})$ is identical to the metric equation of motion Eq. (7) with the substitution $H=H_{\mathrm{eff}}^{\dagger}$. Note however, that the initial conditions are different. For example, a natural choice of initial condition for the metric is $\rho(0)=\mathbbm{1}$, as is discussed in Sec.III and Ref. [46]. On the other hand, a density matrix always take the form $\rho_{\mathrm{nj}}(0)=\frac{1}{2}(\mathbbm{1}+\mathbf{r}\cdot\sigma),|\mathbf{r}|\leq 1$. Note that $\rho_{\mathrm{nj}}(t)$ does not describe a mixed state and is not trace-preserving [50, 51]. In the no-jump limit, the expectation value of an observable is defined as $\langle\hat{O}\rangle_{\text{nj}}=\text{Tr}(\rho_{\mathrm{nj}}\hat{O})/\text{Tr}(\rho_{\mathrm{nj}})$. If the initial state is a pure state, the solution to the no-jump master equation is given by $\rho_{\mathrm{nj}}(t)=\lvert\psi(t)\rangle\langle\psi(t)\lvert$ where $i\dot{\lvert\psi(t)\rangle}=H_{\mathrm{eff}}\lvert\psi(t)\rangle$. Consequently, [35, 36, 37, 38, 27, 39]:
@@ III Metric as a generalization of BQM
In this section, we show that the metric framework is a generalization of the consistent non-Hermitian quantum mechanics encapsulated by BQM. We proceed by first highlighting the equivalence between BQM and a particular similarity transformation, known as pseudo-Hermiticity in the literature [52, 53]. This then motivates the metric framework as a natural generalization of BQM to a dynamical context. For clarity, we consider a general time-independent Hamiltonian $H$ with a real, non-degenerate spectrum and an appropriately normalized biorthogonal basis fulfilling Eqs. (2) and (3).
@@
In the biorthogonal basis, we can write $S$ and $S^{-1}$ as
@@
such that $SS^{-1}=\mathbbm{1}$ by Eqs. (2) and (3). We have $S\lvert n\rangle_{R}=\lvert n\rangle_{L}$ and $S^{-1}\lvert n\rangle_{L}=\lvert n\rangle_{R}$, and so $S\lvert\psi\rangle=\lvert\tilde{\psi}\rangle$ according to Eqs. (4) and (5). Furthermore, for any operator $F$, we have the equivalence
$\langle\psi\lvert SF\lvert\psi\rangle=\langle\tilde{\psi}\lvert F\lvert\psi\rangle$ (16)
and in particular $\langle\psi\lvert S\lvert\psi\rangle=\langle\tilde{\psi}\lvert\psi\rangle=\sum_{n}|c_{n}|^{2}$.
@@
Equation (17) closely resembles the equivalence of the Hilbert spaces $\mathscr{H}$ and $\mathscr{H}_{\rho(t)}$ in the metric framework, in the particular case where the metric is time-independent. This is clear when we identify $\sqrt{S}\rightarrow\eta$, $\tilde{F}\rightarrow H$ and $F\rightarrow h$ in Eqs. (7) and (8). In fact, the matrix $S$, which satisfies Eq. (14), is a time-independent solution to the metric equation of motion Eq. (7) with the initial condition $\rho(0)=S$, such that $\dot{\rho}=0$. It is then evident that the notions in the metric framework are generalizations of the concepts discussed above, namely BQM and pseudo-Hermiticity, to time-dependent cases.
With respect to dynamics, we emphasize that the initial condition for the metric plays an important role in the time evolution of a non-Hermitian system. The metric equation of motion, Eq. (7), guarantees that $\frac{d}{dt}\langle\psi(t)\lvert\rho(t)\lvert\psi(t)\rangle=0$, so $\langle\psi(t)\lvert\rho(t)\lvert\psi(t)\rangle=\langle\psi(0)\lvert\rho(0)\lvert\psi(0)\rangle$. This prompts the choice of initial condition $\rho(0)=\mathbbm{1}$[46], which ensures that $\langle\psi(t)\lvert\rho(t)\lvert\psi(t)\rangle=\langle\psi(0)\lvert\psi(0)\rangle=1$ given a normalized initial state $\lvert\psi(0)\rangle$. Alternatively, the initial condition $\rho(0)=S$ gives rise to a stationary metric, as discussed earlier. With this choice, we would have $\langle\psi(t)\lvert\rho(t)\lvert\psi(t)\rangle=\langle\psi(t)\lvert S\lvert\psi(t)\rangle=\sum_{n}|c_{n}(t)|^{2}$. As discussed in Sec. II.1, this is not necessarily equal to one. Evidently, the dynamical nature of the metric, coupled with an appropriate initial condition, plays an important role in ensuring the subsequent normalization of the states. This is unsurprising, since the metric formalism encompasses a standard Schrödinger-like evolution of a closed quantum system at its core.
@@ IV.1 $\Gamma^{\dagger}\Gamma$ commutes with system Hamiltonian
We consider a dissipative system with system Hamiltonian $H=\omega\sigma_{z}$ and jump operator $\Gamma=2\sqrt{\gamma}\sigma_{-}=\sqrt{\gamma}(\sigma_{x}-i\sigma_{y})$, where $\omega,\gamma>0$. For example, this corresponds physically to the decay process of a two-level system from the excited to the ground state. The effective non-Hermitian Hamiltonian, which appears in the no-jump master equation, is given by $H_{\mathrm{eff}}=(\omega-i\gamma)\sigma_{z}$ up to a constant term proportional to the identity operator. We note that the energy eigenvalues are always complex in this model.
@@
However, for initial conditions that are superpositions of both the ground and excited states, we observe different behaviors. With the initial condition $\lvert\psi(0)\rangle=\frac{1}{\sqrt{2}}(\lvert 0\rangle-\lvert 1\rangle)$, the metric observable oscillates indefinitely as
@@ IV.2 $\Gamma^{\dagger}\Gamma$ does not commute with system Hamiltonian
The effective non-Hermitian Hamiltonian is $H_{\mathrm{eff}}=\omega\sigma_{x}-i\gamma\sigma_{z}$ up to a term proportional to the identity operator. The energy eigenvalues are ±$\pm\sqrt{\omega^{2}-\gamma^{2}}$. There are two parameter regimes: $\mathcal{PT}$-symmetric regime where $\gamma^{2}\leq\omega^{2}$ and a regime where the $\mathcal{PT}$-symmetry is spontaneously broken, $\gamma^{2}>\omega^{2}$. In the regime $\gamma^{2}>\omega^{2}$, the energy eigenvalues are purely imaginary and appear as complex conjugate pairs. The significance of the different parameter regimes stems from the fact that the effective Hamiltonian $H_{\mathrm{eff}}$ now features an exceptional point at $\omega=\gamma$. The solutions obtained by the no-jump approximation and the metric framework display parameter-dependent behaviors: in the $\mathcal{PT}$-broken regime $\omega^{2}<\gamma^{2}$, the solutions tend to a constant value in the limit $t\rightarrow\infty$, whereas in the $\mathcal{PT}$symmetric regime $\omega^{2}>\gamma^{2}$, both methods yield oscillatory solutions. The only qualitative difference is that the metric expectation value displays periodic oscillations, while the oscillations of no-jump expectation value are aperiodic.
@@
To summarize, the full master equation, the no-jump approximation and the metric approaches all display different behaviors in general, despite being model and parameter-specific. For the simpler model in Sec. IV.1, the metric solution is oscillatory in general and does not tend to an asymptotic value in the long-time limit. On the other hand, the qualitative behavior of the no-jump solution resembles the master equation solution. In the long-time limit, the asymptotic value of the no-jump solution coincides with the steady state solution of the master equation solution. The role of the exceptional points in the dynamics is evident in the more complicated model discussed in Sec. IV.2. In the parameter regime where $\mathcal{PT}$-symmetry is preserved, there are common qualitative features between the metric and no-jump expectation values, which is oscillatory behavior. They only tend to an asymptotic value, resembling the steady state solution, in the parameter regime where $\mathcal{PT}$-symmetry is spontaneously broken. These asymptotic values generally do not coincide with the steady state solution of the master equation, though the no-jump approximation reproduces the qualitative behavior of the master equation albeit only in the $\mathcal{PT}$-broken regime. We also found that the asymptotic values of the metric expectation values are sensitive to initial conditions. This sensitivity can be attributed to Schrödinger evolution with an intrinsically non-Hermitian Hamiltonian, in contrast to the dissipative dynamics described by the full master equation. These observations motivates further study of whether there are a class of initial states where the metric formalisms leads to observables which mimic steady state values. Our findings thus highlight the importance of considering the physical origin of non-Hermiticity when studying the dynamics of non-Hermitian systems.

### v2

@@ I Introduction
Dynamics of non-Hermitian systems is a new frontier of exploration which harbors a lot of potential to discover new physical paradigms. A plethora of novel phenomena, such as dynamically induced topological phases, exponentially-enhanced quantum sensing [26] and the divergence of Lieb-Robinson bounds [8, 27] have been reported recently. Despite its increasing popularity, there is a lack of consensus in the mathematical formulation of non-Hermitian quantum mechanics in the community. As non-Hermitian time evolution is intrinsically non-unitary, different methods have been adopted to obtain a normalized physical observable in non-Hermitian systems. Among them is biorthogonal quantum mechanics (BQM) [28], which is a comprehensive formulation of non-Hermitian quantum mechanics consistent with its probabilistic interpretation. BQM is widely adopted within the condensed matter community [29, 30, 31, 32] and mainly applied to study systems featuring a time-independent Hamiltonian. The so-called metric framework  [33, 34, 35, 36] provides a generalization of BQM to the case with time-dependent Hamiltonians. In this framework, one considers a modified Hilbert space endowed by a time-dependent metric which appears as a weight factor in the inner product. The equation of motion of the metric incorporates the dynamics induced by both the Hamiltonian and its Hermitian conjugate. Both of these methods generate an exact time evolution of the non-Hermitian system, and are best suited when the non-Hermiticity of the system is realized in a fundamental manner instead of arising as an approximation. For example, this can be realized in NV centres via Naimark dilation [12].
On the other hand, a lot of the available literature on non-Hermitian dynamics [37, 38, 39, 40, 27, 41] skirts the important issue of non-conserved norms during time evolution and simply divides physical quantities by the time-dependent norm of the time-evolved state. This approach is ubiquitous in the open quantum systems community and is motivated by the quantum trajectory theory [42, 43, 44, 45, 46]. Quantum trajectories are used to find the time evolution of open quantum systems [47], for example when the direct solution of Lindblad master equation is computationally costly. In certain scenarios such as the detection of photons emitted from a cavity, the evolution of the system can be approximated by that with an effective non-Hermitian Hamiltonian interrupted by quantum jumps at random times [42]. By post-selecting quantum trajectories which did not undergo quantum jumps, a non-Hermitian evolution can be realized and the time-evolved state has to be explicitly normalized by hand at every time step [37].
@@ II.1 Biorthogonal quantum mechanics (BQM)
We start with a brief summary of BQM [28] for a general static non-Hermitian Hamiltonian $H$ with a non-degenerate spectrum. We assume a finite-dimensional Hilbert space throughout the paper. The inifinite-dimensional case has been discussed in Refs. [28, 51]. The left and right eigenstates of $H$ are defined as
@@
as expectation values of observables in the biorthogonal framework are defined as $\langle\hat{F}\rangle=\tilde{\langle\psi\lvert}\hat{F}\lvert\psi\rangle$. In particular, setting $\hat{F}=\mathbbm{1}$, we see $\langle\tilde{\psi}\lvert\psi\rangle=\sum_{n}|c_{n}|^{2}$=1.
@@ II.2 Metric framework
Time evolution in the Hilbert spaces $\mathscr{H}_{\rho(t)}$   and $\mathscr{H}$  is generated by the respective Hamiltonians, $H(t)$ and $h(t)$, via the time-dependent Schrödinger equation
@@
where the states are related by $|\Psi(t)\rangle=\eta(t)|\psi(t)\rangle$. The construction of $\rho(t)$ guarantees that $h(t)$ is Hermitian, such that the unitarity of the time evolution is restored. In particular, Eq. (7) guarantees that $\langle\psi(t)|\rho(t)|\psi(t)\rangle=\langle\psi(0)\lvert\rho(0)\lvert\psi(0)\rangle$ at all times $t$. A natural choice of initial condition is then $\rho(0)=\mathbbm{1}$[50], which ensures that $\langle\psi(t)\lvert\rho(t)\lvert\psi(t)\rangle=\langle\Psi(t)|\Psi(t)\rangle=1$ given a normalized initial state $\lvert\psi(0)\rangle=\lvert\Psi(0)\rangle$.
@@ II.3 Open quantum systems
and neglecting the term $\Gamma\rho_{\mathrm{me}}\Gamma^{\dagger}$, which is known as the recycling or quantum jump term. Curiously, the approximate master equation in the no-jump limit, $\dot{\rho}_{\mathrm{nj}}=-i(H_{\mathrm{eff}}\rho_{\mathrm{nj}}-\rho_{\mathrm{nj}}H_{\mathrm{eff}}^{\dagger})$ is identical to the metric equation of motion Eq. (7) with the substitution $H=H_{\mathrm{eff}}^{\dagger}$. Note however, that the initial conditions are different. For example, a natural choice of initial condition for the metric is $\rho(0)=\mathbbm{1}$, as is discussed in Sec.III and Ref. [50]. On the other hand, the initial state $\rho_{\mathrm{nj}}(0)$ is usually taken to be a density matrix. For example, the density matrix of a qubit always takes the form $\rho_{\mathrm{nj}}(0)=\frac{1}{2}(\mathbbm{1}+\mathbf{r}\cdot\sigma),|\mathbf{r}|\leq 1$. Note that the time-evolved $\rho_{\mathrm{nj}}(t)$ does not describe a mixed state and is not trace-preserving [55, 56]. In the no-jump limit, the expectation value of an observable is defined as $\langle\hat{O}\rangle_{\text{nj}}=\text{Tr}(\rho_{\mathrm{nj}}\hat{O})/\text{Tr}(\rho_{\mathrm{nj}})$. If the initial state is a pure state, the solution to the no-jump master equation is given by $\rho_{\mathrm{nj}}(t)=\lvert\psi(t)\rangle\langle\psi(t)\lvert$ where $i\dot{\lvert\psi(t)\rangle}=H_{\mathrm{eff}}\lvert\psi(t)\rangle$. Consequently, [37, 38, 39, 40, 27, 41]:
@@ III Metric as a generalization of BQM
In this section, we show explicitly that the metric framework is a generalization of the consistent non-Hermitian quantum mechanics encapsulated by BQM, introduced in Sec. II.1. We first summarize the findings of Refs. [28, 57], which show the equivalence between BQM and a particular similarity transformation known as pseudo-Hermiticity [58, 59, 57]. We then illustrate how the time-dependent metric extends the notion of pseudo-Hermiticity to dynamical cases [34]. This motivates the metric framework as a natural generalization of BQM to a dynamical context.
For clarity, we consider a general time-independent Hamiltonian $H$ with a real, non-degenerate spectrum and an appropriately normalized biorthogonal basis fulfilling Eqs. (2) and (3). For a Hamiltonian with degenerate spectrum, the left and right eigenstates may not form complete sets, in which case the similarity transformation given by Eq. (15) does not exist [28].
@@
In the biorthogonal basis, we can write $S$ and $S^{-1}$ as [28, 57]
@@
such that $SS^{-1}=\mathbbm{1}$ by Eqs. (2) and (3). We have $S\lvert n\rangle_{R}=\lvert n\rangle_{L}$ and $S^{-1}\lvert n\rangle_{L}=\lvert n\rangle_{R}$, and so $S\lvert\psi\rangle=\lvert\tilde{\psi}\rangle$ according to Eqs. (4) and (5). Furthermore, for any operator $F$, we have the equivalence [28]
$\langle\psi\lvert SF\lvert\psi\rangle=\langle\tilde{\psi}\lvert F\lvert\psi\rangle.$ (16)
This shows the equivalence of observables obtained using BQM and by considering the pseudo-Hermiticity transformation (15). We also note that $\langle\psi\lvert S\lvert\psi\rangle=\langle\tilde{\psi}\lvert\psi\rangle=1$.
@@
Equation (17) closely resembles the equivalence of the Hilbert spaces $\mathscr{H}$ and $\mathscr{H}_{\rho(t)}$ in the metric framework, in the particular case where the metric is time-independent. This is clear when we identify $\sqrt{S}\rightarrow\eta$, $\tilde{F}\rightarrow H$ and $F\rightarrow h$ in Eqs. (7) and (8). In fact, the matrix $S$, which satisfies Eq. (14), is a time-independent solution to the metric equation of motion Eq. (7) with the initial condition $\rho(0)=S$, such that $\dot{\rho}=0$. A caveat is that not all non-Hermitian Hamiltonians can be mapped to its Hermitian conjugate by a similarity transformation. The existence of such a mapping requires the eigenvalues to be real and non-degenerate [52], so a time-independent solution $S$ to Eq. (7) does not generically exist. However, even when a stationary solution is not permissible, a dynamical solution $\rho(t)$ can be found. It is then evident that the notions in the metric framework are generalizations of the concepts discussed above, namely BQM and pseudo-Hermiticity, to time-dependent cases.
@@ IV.1 $\Gamma^{\dagger}\Gamma$ commutes with system Hamiltonian
We consider a dissipative system with system Hamiltonian $H=\omega\sigma_{z}$ and jump operator $\Gamma=2\sqrt{\gamma}\sigma_{-}=\sqrt{\gamma}(\sigma_{x}-i\sigma_{y})$, where $\omega,\gamma>0$. For example, this corresponds physically to the decay process of a two-level system from the excited to the ground state. The effective non-Hermitian Hamiltonian, which appears in the no-jump master equation, is given by $H_{\mathrm{eff}}=(\omega-i\gamma)\sigma_{z}$ up to a constant $i\gamma\mathbbm{1}$ term. This term contributes to some factors in the time-evolved state and the metric. However, no matter which method we use (metric or division by norm), the factor cancels out when computing physical observables and therefore can be neglected. We note that the energy eigenvalues are always complex in this model.
@@
However, for initial conditions that are superpositions of both the ground and excited states, we observe different behaviors. With the initial condition $\lvert\psi(0)\rangle=\frac{1}{\sqrt{2}}(\lvert 0\rangle-\lvert 1\rangle)$, the metric observable oscillates indefinitely:
@@ IV.2 $\Gamma^{\dagger}\Gamma$ does not commute with system Hamiltonian
The effective non-Hermitian Hamiltonian is $H_{\mathrm{eff}}=\omega\sigma_{x}-i\gamma\sigma_{z}$ up to a term proportional to the identity operator. The exact solutions of Eq. (9) and (8) are given in Appendix A. The energy eigenvalues are ±$\pm\sqrt{\omega^{2}-\gamma^{2}}$. There are two parameter regimes: $\mathcal{PT}$-symmetric regime where $\gamma^{2}\leq\omega^{2}$ and a regime where the $\mathcal{PT}$-symmetry is spontaneously broken, $\gamma^{2}>\omega^{2}$. In the regime $\gamma^{2}>\omega^{2}$, the energy eigenvalues are purely imaginary and appear as complex conjugate pairs. The significance of the different parameter regimes stems from the fact that the effective Hamiltonian $H_{\mathrm{eff}}$ now features an exceptional point at $\omega=\gamma$. The solutions obtained by the no-jump approximation and the metric framework display parameter-dependent behaviors: in the $\mathcal{PT}$-broken regime $\omega^{2}<\gamma^{2}$ shown by Figs. 1(b) and (d), the solutions tend to a constant value in the limit $t\rightarrow\infty$, whereas in the $\mathcal{PT}$symmetric regime $\omega^{2}>\gamma^{2}$ shown by Figs. 1(a) and (c), both methods yield oscillatory solutions. The only qualitative difference is that the metric expectation value displays periodic oscillations, while the oscillations of no-jump expectation value are aperiodic.
@@
To summarize, the full master equation, the no-jump approximation and the metric approaches all display different behaviors in general, despite being model and parameter-specific. For the simpler model in Sec. IV.1, the metric solution is oscillatory in general and does not tend to an asymptotic value in the long-time limit. On the other hand, the qualitative behavior of the no-jump solution resembles the master equation solution. In the long-time limit, the asymptotic value of the no-jump solution coincides with the steady state solution of the master equation solution. The role of the exceptional points in the dynamics is evident in the more complicated model discussed in Sec. IV.2. In the parameter regime where $\mathcal{PT}$-symmetry is preserved, there are common qualitative features between the metric and no-jump expectation values, which is oscillatory behavior. They only tend to an asymptotic value, resembling the steady state solution, in the parameter regime where $\mathcal{PT}$-symmetry is spontaneously broken. These asymptotic values generally do not coincide with the steady state solution of the master equation, though the no-jump approximation reproduces the qualitative behavior of the master equation albeit only in the $\mathcal{PT}$-broken regime. We also found that the asymptotic values of the metric expectation values are sensitive to initial conditions. This sensitivity can be attributed to Schrödinger evolution with an intrinsically non-Hermitian Hamiltonian, in contrast to the dissipative dynamics described by the full master equation. These observations motivates further study of whether there are a class of initial states where the metric formalisms leads to observables which mimic steady state values. Our findings thus highlight the importance of considering the physical origin of the putative non-Hermiticity, since this dictates the choice of methodology so as to obtain the correct physical predictions. Namely, when non-Hermiticity is considered in the no-jump limit of a dissipative system, BQM/metric should not be used but instead the norm-method should be. This also demonstrates that a self-consistent non-Hermitian quantum system encapsulated by the metric formalism might harbor very different paradigms of physics as compared to the standard no-jump limit of dissipative systems.